OpenStudy (anonymous):

1.Create a quadratic function, f(x), in vertex form. The a should be between 4 and –4, the h will be your birth month, and the k will be your birth day. Write your equation below. 2.Using complete sentences, explain how to convert your birthday function into standard form. 3.Graph your function. Include your graph below. 4.Using complete sentences, explain how to find the average rate of change for f(x) from x = 4 to x = 7.

OpenStudy (anonymous):

mine is January 3rd 1998

OpenStudy (triciaal):

h = Month = January =1 k = day= 3 vertex form f(x) = a(x-h)^2 + k you need to choose your value for a The a should be between 4 and –4, substitute these values into the f(x), the equation

OpenStudy (triciaal):

explain how to convert your birthday function into standard form. standard form f(x) = ax^2 + bx + c = 0 f(x) = 0

OpenStudy (triciaal):

3.Graph your function. Include your graph below. find the 2 values of x when f(x), y = 0 when x = 0 y = c plot the points (x1, 0), (0, C), (x2, 0) where x1 and x2 are the values of x when y = 0

OpenStudy (triciaal):

explain how to find the average rate of change for f(x) from x = 4 to x = 7. the rate of change is the slope

OpenStudy (anonymous):

wait okay so did u just answer my problem? like on step 1?

OpenStudy (anonymous):

im still confused. can we do this together step by step?